OpenStudy (arihangdu):

A radioactive nucleus alpha decays to yield a sodium-24 nucleus in 14.8 hours. What was the identity of the original nucleus? Show the nuclear equation.

OpenStudy (cuanchi):

an alpha particle is equal to a He nucleus. It has two protons and 2 neutrons, then the mass of the nucleotide is \[_{2}^{4}He =_{2}^{4}\alpha \] \[_{?}^{?}X \rightarrow _{11}^{24}Na + _{2}^{4}He \]

OpenStudy (cuanchi):

you have to find the reactant in this equation, the mass and the number of protons has to be the same in both sides of the equation you add 24+4= 28 and 11+2= 13 the 13 atomic number is telling you what element it is and the 28 atomic mass is telling you what is the mass of the isotope of that element

OpenStudy (arihangdu):

Thanks you

OpenStudy (arihangdu):

X aluminum right?

OpenStudy (cuanchi):

yes, and the mass?

OpenStudy (arihangdu):

But in the period table it is Si with mass 28 right or

OpenStudy (arihangdu):

Al has only 26.982 in my table ...

OpenStudy (cuanchi):

No dear!!! that is the average mass proportional to the abundance of the different isotopes. In this problem you have only one of the isotopes of the Al. The mass is equal to the sum of the mass of the nucleoids In your case is \[_{13}^{28}Al\] the isotope of the Al that has a mass of 28 You are right Si can have also some isotope with a mass of 28 or the average mass is 28 but what tell you what element it is is the number of protons

OpenStudy (arihangdu):

Oke thank you my friend appreciate your help with good explanation.you are great.

OpenStudy (cuanchi):

Al has many isotopes https://en.wikipedia.org/wiki/Isotopes_of_aluminium

OpenStudy (arihangdu):

Okay thanks bye I don't want to kill more your time. Thanks you very much.byeeee great help for me